OpenStudy (anonymous):

The slope of a line is 12 and the point (1, 10) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

y-10 = 12 ----- x-1 y-10=12(x-1) y-10=12x-12 y=12x-2

OpenStudy (anonymous):

you mean +2

OpenStudy (anonymous):

no its -2 coz (-12+10=-2)